# Personalized Path Recourse for Reinforcement Learning Agents

### v2

@@ Abstract
This paper introduces Personalized Path Recourse, a novel method that generates recourse paths for a reinforcement learning agent. The goal is to edit a given path of actions to achieve desired goals (e.g., better outcomes compared to the agent’s original path) while ensuring a high similarity to the agent’s original paths and being personalized to the agent. Personalization refers to the extent to which the new path is tailored to the agent’s observed behavior patterns from their policy function. We train a personalized recourse agent to generate such personalized paths, which are obtained using reward functions that consider the goal, similarity, and personalization. The proposed method is applicable to both reinforcement learning and supervised learning settings for correcting or improving sequences of actions or sequences of data to achieve a pre-determined goal. The method is evaluated in various settings. Experiments show that our model not only recourses for a better outcome but also adapts to different agents’ behavior.
Machine Learning, ICML
@@ 1 Introduction
Imagine John, a 30-year-old accountant earning an annual income of \$50,000, who wishes he had pursued a different life path that could have led to a higher income of over \$100,000. This situation raises a fundamental question: how can we generate alternative paths of actions that might have resulted in a better outcome?
@@
There are many alternative paths that John could have taken to earn more than \$100,000, such as pursuing a computer science degree when he was in college or accepting a different job offer a couple of years ago, or, more radically, starting sports training when he was young to become a professional athlete. However, which path is the most suitable for John? First, when achieving the desired goal, we prefer smaller changes to John’s life path. Thus the new path needs to bear as much similarity as possible to the original path. Second, we want the new path to be personalized, or tailored, to John. This means that the new path of actions should respect John’s behavior patterns and preferences. For instance, if John is interested in mathematics (as also reflected by his current job as an accountant) but doesn’t show any talent in sports, recommending a path that leads him to become a data scientist would be more suitable than suggesting him to become an athlete. We call such a new path that considers the alignment with an agent’s own behavior pattern and similarity to the agent’s original path while achieving a desired goal personalized recourse.
Both similarity and personalization hold significant importance in real-world applications. Similarity is important as users tend to be more receptive to recommendations that align closely with their initial preferences. Consider the case of a driving route. A route that closely resembles a driver’s usual path is often preferred, as the driver is already familiar with its surroundings, traffic conditions, etc. In the context of coaching, individuals are likely to learn more effectively when provided with feedback based on their past performance rather than being advised to adopt an entirely different strategy. Personalization is also necessary since a recommended recourse should respect how an individual would intrinsically take actions in different scenarios. Note that personalization is different from similarity: similarity is evaluated at the path level, while personalization is evaluated at the policy level, i.e., the likelihood of taking different actions in a given state.
@@
This personalized recourse is different from but related to counterfactual explanations or algorithmic recourse, which have predominantly focused on supervised learning (Verma et al., 2020; Pawlowski et al., 2020; Tsirtsis & Gomez Rodriguez, 2020; Goyal et al., 2019): how to change one or some of the features of an instance $\mathbf{X}$ to get a different outcome $Y$? Out of the few relevant works we found for reinforcement learning, counterfactual reasoning is used for the purpose of explaining why an action is chosen for a given state $\mathbf{s}$ over the other actions, by generating a counterfactual state $\mathbf{s}^{\prime}$ that could lead to a different action (Olson et al., 2019; Madumal et al., 2020; Gajcin & Dusparic, 2023). However, these methods do not deviate from similar approaches in supervised learning because the explanations are on the state level and still focus on one decision (action) at a time (Gajcin & Dusparic, 2022). Frost et al. (2022) extends the explanation scope from states to paths, but still for the purpose of explanations: they create a set of informative paths of the agent’s behavior under a test-time state distribution but these paths do not necessarily lead to a better outcome.
@@
None of the previously mentioned papers solve our problem, which focuses on “recourse” rather than “explanation”, with the goal of creating an entire path of actions that can lead an agent to a desired goal (e.g., better outcome), while being tailored to the agent.
We work with sequence data that can be modeled by Markov Decision Processes (MDP) (Puterman, 2014). In MDP, an agent can be trained to learn a policy using reinforcement learning. Given a path of states and actions from a focal agent $A$, our method, Personalized Path Recourse (PPR), trains a new policy to generate recourse paths of actions that satisfy similarity and personalization requirements while leading to a better outcome. Our framework also works for sequence data in other settings, such as generative AI, or supervised learning settings. In those settings, an “agent” is the underlying data-generating process of the sequences. Then the problem is equivalent to how to modify the sequence in order to achieve a certain goal (e.g., flipping the label) while making sure the generated sequence is plausible and in distribution.
To train such a personalized recourse agent, we devise a reward function that considers three aspects, similarity to the original path, personalization (i.e., how much it is tailored to agent $A$), and goal (e.g., a better outcome). We also design an efficient exploration function to make sure the algorithm works on large state space and action space.
PPR is evaluated in different settings (reinforcement learning and supervised learning). For comparison, we adapted and modified a few methods to our setting (we did not find any existing methods that can directly solve our problem) and experiments show that PPR achieves more personalized paths while achieving a pre-determined goal. Given different agents, PPR is able to personalize the recourse path to adapt to the specific styles of different agents.
@@ 2.1 Algorithmic Recourse
The problem PPR solves is closely related to the goal of algorithmic recourse (Karimi et al., 2021), the process by which one can change an unfavorable outcome or decision made by an algorithm by altering certain input variables within their control. Algorithmic recourse can be considered as a special case of counterfactual explanations (Wachter et al., 2017; Dandl et al., 2020; Thiagarajan et al., 2021). In this line of research, the input instance is represented by a fixed-length vector (Poyiadzi et al., 2020; Verma et al., 2022; De Toni et al., 2023; Wang et al., 2021; Van Looveren et al., 2021; Sulem et al., 2022; Karlsson et al., 2020; Hsieh et al., 2021; Ates et al., 2021). Some work considers the feasibility and costs of the changes made to the features (Ustun et al., 2019; Ross et al., 2021; Joshi et al., 2019; Upadhyay et al., 2021; Harris et al., 2022; Karimi et al., 2020), and also personalization or user preference (Yetukuri et al., 2022; De Toni et al., ). However, these existing works have solely focused on supervised learning in tabular data settings, where our method can work with both structured and unstructured data in a reinforcement learning setting.
@@ 2.2 Counterfactual Explanations for Reinforcement Learning
The majority of the work on counterfactual explanations is for supervised learning environments. Only a small number of research have been done for problems in reinforcement learning settings, and they are mostly focused on explanations for one action at a time. Various methods (Olson et al., 2019; Gajcin & Dusparic, 2023; Madumal et al., 2020) have been proposed to generate counterfactual states that could lead to a different action when explaining an agent’s current action for a given state. All these works provide counterfactual explanations for one state and action at a time. Frost et al. (2022) provides explanations at the path level, which trains an exploration policy to generate test time paths to explain how the agent behaves in unseen states following training. However, all these previously mentioned papers aim to explain rather than provide a complete solution, namely a new path, towards a better outcome.
Tsirtsis et al. (2021) is the most relevant to our task. The paper focuses on sequential decision-making under uncertainty and generates counterfactual explanations that specify an alternative sequence of actions differing in at most $k$ actions from the observed sequence that could have led to a better outcome. However, by employing dynamic programming and iterating through all the possible states and actions, the method only solves tasks with low-dimensional states and action spaces, as later shown in the experiments. Additionally, only comparing which actions are changed is not meaningful in many applications because the entire routes are already different after taking different actions, even if the following actions are the same. Finally, this method also does not consider personalization to an agent when looking for a counterfactual sequence of actions. This method will be later used as a baseline for some simple tasks with low complexity.
@@ 2.3 Alternative Path Planning
In addition, there is a set of works that find alternative policies for an agent by modifying its learning process (Racanière et al., 2017; Zahavy et al., 2021; Andrychowicz et al., 2017). However, those works aim to discover either more diverse or potentially better policies. While their ideas can help improve the agent’s training process, none of them could help solve finding an agent’s personalized recourse path.
@@ 3.1 Background and Notations
An MDP is a tuple M $\triangleq(\mathcal{S},\mathcal{A},P,R,\lambda$), where $\mathcal{S}=\{\mathbf{s}_{1},...,\mathbf{s}_{|\mathcal{S}|}\}$ and $\mathcal{A}=\{a_{1},...,a_{|\mathcal{A}|}\}$ are the set of states and actions an agent $A$ can make in the environment. Here, $|\mathcal{S}|$, $|\mathcal{A}|$ are the cardinalities of set $\mathcal{S}$ and $\mathcal{A}$ respectively. $\lambda\in$ [0, 1) is the discount factor. At time $t$, the agent at state $\mathbf{s}_{t}$ takes an action $a_{t}$ and move to state $\mathbf{s}_{t+1}$ with a transition probability $P(\mathbf{s}_{t+1}|\mathbf{s}_{t},a_{t})$ and receives a reward, $r_{t}$, which is determined by a reward function $R(\mathbf{s}_{t},a_{t})$. Behavior of the agent is defined by a policy function $\mathcal{P}_{A}(\mathbf{s}_{t},a_{t})$, which defines the probability that the agent $A$ will choose action $a_{t}$ from state $\mathbf{s}_{t}$ at time $t$. An episode is defined as a sequence of interactions with the environment made by the agent from the beginning to the end. We denote $\tau=<\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{T}>$ as a route or path, i.e., a sequence of states the agent $A$ was in this episode. In general, the goal of RL is to train the agent to learn an optimal, or nearly-optimal, policy $\pi(\mathbf{s},a)$ to help it maximize the reward from the environment.
@@
There are many algorithms for training a reinforcement learning agent. In this paper, we use deep Q-networks (DQN) to train a recourse policy. In a DQN algorithm, first, the agent explores the environment in many episodes. For each episode, at each time step $t$, we record the agent’s current state $\mathbf{s}_{t}$, action $a_{t}$, the next state $\mathbf{s}_{t+1}$ and the reward $r_{t}$. Those tuples $(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})$ are stored in an experience replay $B$. DQN trains a deep neural network parameterized by $\theta$, $Q_{\theta}(\mathbf{s}_{t},a_{t})$, that helps the agent to maximize the reward from the environment. To facilitate the training, a target network is created, $Q_{\theta^{\prime}}(\mathbf{s}_{t},a_{t})$, parameterized by $\theta^{\prime}$. $\theta^{\prime}$ is updated every few time steps by setting $\theta^{\prime}=\theta$. Creating a lag in updating the target network helps the training of the original network $Q_{\theta}$ to become more stable. With experience replay and the target network introduced, the DQN minimizes the following loss function:
$\begin{split}L(\theta)=\operatorname{\mathbb{E}}_{(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})\sim\mathcal{V}(B)}[(r_{t}+\\ \gamma*\max_{a_{t}}Q_{\theta^{\prime}}(\mathbf{s}_{t+1},a_{t})-Q_{\theta}(\mathbf{s}_{t},a_{t}))]\end{split}$ (1)
Here, mini batches of tuples $(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})$ are sampled from the replay buffer $B$ with sample strategy $\mathcal{V}$, such as uniform distribution. Note that other extensions from DQN such as prioritized replay buffer (Schaul et al., 2015) or dueling networks (Wang et al., 2016) can also be applied.
@@ 3.2 Problem Formulation
Goal The goal $\mathcal{G}$ could represent different constraints or desired outcomes, depending on the task. For example, it could require agent A to arrive at a particular destination state $\mathbf{s}_{d}$; or it could require the overall reward to be larger than some pre-defined threshold (e.g., a driver reaching the final destination faster); or in the context of sequence data in a supervised learning setting, it could require the sequence to be classified as positive. The goal can be formulated as a continuous outcome, the higher (lower) the better, or a binary outcome, whether certain criteria are met.
Similarity This property requires the $\tau_{r}$ to be as similar to the original route $\tau_{0}$ as possible. For example, the new route $\tau_{r}$ recommended to the driver should have many overlapped states with the original route $\tau_{0}$ picked by the driver.
Personalization This property requires the new route $\tau_{r}$ to have a high likelihood of being generated by the agent $A$. For example, if $\tau_{r}$ is a new route recommended to a driver, with the goal to reach some destination as fast as possible, we would recommend a highway to an experienced driver and a local route to a new driver. In other words, the recommended new route $\tau_{r}$ needs to respect agent $A$’s behavior. In practice, the agent’s behavior is usually learned from a separate data stream or provided by an external system, which could be in many different formats. Some examples of this could be the language models such as ChatGPT, which users can only use as-a-service; or streaming companies usually have a dedicated system/database to query users’ choices when being given a set of recommendations. In our method, we use policy function $\mathcal{P}_{A}(.)$ to represent agent A’s behavior. This approach could help impose generalizability, because $\mathcal{P}_{A}(.)$ is usually already pre-trained/provided (language models, RL agent’s policies, etc.), or could be trained from agent A’s historical data using various state-of-the-art techniques.
Note that personalization is different from similarity. Personalization encourages the agent to follow a policy $\mathcal{P}_{A}$, which represents the expected behavior of the agent, i.e., what action/choice the agent will take given a state/situation. Therefore, personalization evaluates how much the generated route respects the distribution of behavior of the agent. On the other hand, the similarity is evaluated at the route level, which is specific to a given original route $\tau_{0}$. For example, an agent may have equal probabilities of taking the right route and left route from a crossroad. Suppose the original route $\tau_{0}$ goes right. Then the left route has a low similarity to $\tau_{0}$ but a high personalization score because the agent is equally likely to go left.

### v3

@@ Abstract
This paper introduces Personalized Path Recourse, a novel method that generates recourse paths for a path from an agent. The objective is to edit a given path of actions to achieve desired goals (e.g., better outcomes compared to the agent’s original paths of actions) while ensuring the new path is personalized to the agent. Personalization refers to the extent to which the new path is similar to the original path and tailored to the agent’s observed behavior patterns from their policy function. We train a personalized recourse agent to generate such personalized paths, which are obtained using reward functions that consider the goal and personalization. The proposed method is applicable to both reinforcement learning and supervised learning settings for correcting or improving sequences of actions or sequences of data to achieve a pre-determined goal. The method is evaluated in various settings and demonstrates promising results.
@@ 1 Introduction
Imagine John, a 30-year-old accountant earning an annual income of \$50,000, who wishes he had pursued a different life path that could have led to a higher income of over \$100,000. This situation raises a fundamental question: in a multi-step decision-making setting, given a current path that is considered inferior, how can we generate alternative paths of actions that might have resulted in a better outcome?
@@
There are many alternative paths that John could have taken to earn more than \$100,000, such as pursuing a computer science degree when he was in college or accepting a different job offer a couple of years ago, or, more radically, starting sports training when he was young to become a professional athlete. However, with the goal of making more money, which path is the most suitable for John? The problem is not just to find the best path for achieving the goal, the problem is to find the best path for John to achieve the goal. In other words, the new path needs to be personalized for John.
We define the meaning of personalization in the context of multi-step decision making. First, people generally prefer options that aren’t too different from their initial choices, as these familiar options tend to be more comfortable and less risky. For example, a driver is likely to prefer a path similar to their usual one, as they are familiar with the surroundings and traffic conditions. Similarly, in a coaching context, individuals are more receptive to feedback based on their past performance rather than being advised to adopt an entirely new strategy. This means when achieving the desired goal, we prefer smaller changes to John’s original path. Thus the new path needs to bear as much similarity as possible to the original path. We call this path-level personalization.
Another aspect of personalization we consider the behaviors or policies of individuals. Returning to John’s career example, a new path should be tailored specifically to him, respecting his behavior patterns and preferences. For instance, if John enjoys mathematics (as reflected in his current job as an accountant) but lacks athletic talent, recommending a path that leads him to become a data scientist would be more appropriate than suggesting he become an athlete. This approach follows John’s intrinsic policy, which is shaped by his education history and behavior over time. We call this policy-level personalization.
Therefore, the path-level personalization encourages following the original path as much as possible and the policy-level personalization encourages following the original policy function as much as possible. We call the new path that could help John reach his goal and is personalized for him at both path and policy level personalized recourse.
@@
This personalized recourse is different from but related to counterfactual explanations or algorithmic recourse, which have predominantly focused on supervised learning (Verma et al. 2020; Pawlowski, Coelho de Castro, and Glocker 2020; Tsirtsis and Gomez Rodriguez 2020; Goyal et al. 2019): how to change one or some of the features of an instance $\mathbf{X}$ to get a different outcome $Y$? Out of the few relevant works we found for reinforcement learning, counterfactual reasoning is used for the purpose of explaining why an action is chosen for a given state $\mathbf{s}$ over the other actions, by generating a counterfactual state $\mathbf{s}^{\prime}$ that could lead to a different action (Olson et al. 2019; Madumal et al. 2020; Gajcin and Dusparic 2023). However, these methods do not deviate from similar approaches in supervised learning because the explanations are on the state level and still focus on one decision (action) at a time (Gajcin and Dusparic 2022). Frost et al. (2022) extends the explanation scope from states to paths, but still for the purpose of explanations: they create a set of informative paths of the agent’s behavior under a test-time state distribution but these paths do not necessarily lead to a better outcome.
@@
None of the previously mentioned papers solve our problem, which focuses on “recourse” rather than “explanation”, with the goal of creating an entire path of actions that can lead an agent to a desired goal (e.g., better outcome), while being tailored to the agent. In addition, none of the above work considers the policy level personalization and only studies counterfactuals at the path level.
In practical applications, policy-level personalization is crucial, which ensures that any recommended recourse respects a given policy, describing how an individual would naturally act in different situations. For instance, an experienced driver might prefer the highway because it is faster, while a new driver might choose a local path because it feels safer. These concepts apply broadly: a gamer might want to find a new way to beat a game with a different play style, or a writer might wish to paraphrase a sentence to convey a different sentiment while maintaining a specific genre, such as scientific or fantasy. Nowadays, policies are readily available through pre-trained models or can be easily learned from large datasets. Some examples of this could be the pre-trained language models from a certain corpus, or dedicated system/database from streaming companies to query users’ choices when being given a set of recommendations. In the gaming industry, the policy can also be trained from the data collected from players over time.
In this paper, we introduce a method called Personalized Path Recourse (PPR), which works with sequence data modeled by Markov Decision Processes (MDP) (Puterman 2014). Our approach trains a new policy to generate recourse paths that pursue both path-level and policy-level personalization while leading to a better outcome. To train a personalized recourse agent, we design a reward function that takes into account three critical aspects: similarity to the original path, personalization to the agent’s behavior (as reflected in their policy), and the goal outcome. Our framework is versatile, applicable to various settings, including reinforcement learning and supervised learning, where the challenge is to modify a sequence to achieve a certain goal while ensuring the generated sequence remains plausible. The method ensures scalability in large state and action spaces.
We evaluate PPR in different settings, comparing it against adapted and modified methods, as no existing methods directly address our problem. Our experiments demonstrate that PPR achieves more personalized paths while attaining the desired goal, effectively adapting to the unique styles of different agents.
@@ 2.1 Algorithmic Recourse
The problem PPR solves is closely related to the goal of algorithmic recourse (Karimi, Schölkopf, and Valera 2021), the process by which one can change an unfavorable outcome or decision made by an algorithm by altering certain input variables within their control. Algorithmic recourse can be considered as a special case of counterfactual explanations (Wachter, Mittelstadt, and Russell 2017; Dandl et al. 2020; Thiagarajan et al. 2021). In this line of research, the input instance is represented by a fixed-length vector (Poyiadzi et al. 2020; Verma, Hines, and Dickerson 2022; De Toni, Lepri, and Passerini 2023; Wang et al. 2021; Van Looveren et al. 2021; Sulem et al. 2022; Karlsson et al. 2020; Hsieh, Moreira, and Ouyang 2021; Ates et al. 2021). Some work considers the feasibility and costs of the changes made to the features (Ustun, Spangher, and Liu 2019; Ross, Lakkaraju, and Bastani 2021; Joshi et al. 2019; Upadhyay, Joshi, and Lakkaraju 2021; Harris et al. 2022; Karimi et al. 2020), and also personalization or user preference (Yetukuri, Hardy, and Liu 2022; De Toni, Lepri, and Passerini 2023). However, these existing works have solely focused on supervised learning in tabular data settings, where our method can work with both structured and unstructured data in a reinforcement learning setting.
@@ 2.2 Counterfactual Explanations for Reinforcement Learning
The stream of work that is closely relevant to ours is counterfactual explanations, though they only focus on the path-level similarity. The majority of the work on counterfactual explanations is for supervised learning environments. Only a small number of research have been done for problems in reinforcement learning settings, and they are mostly focused on explanations for one action at a time. Various methods (Olson et al. 2019; Gajcin and Dusparic 2023; Madumal et al. 2020) have been proposed to generate counterfactual states that could lead to a different action when explaining an agent’s current action for a given state. All these works provide counterfactual explanations for one state and action at a time. Frost et al. (2022) provides explanations at the path level, which trains an exploration policy to generate test time paths to explain how the agent behaves in unseen states following training. However, all these previously mentioned papers aim to explain rather than provide a complete solution, namely a new path, towards a better outcome.
Tsirtsis, De, and Rodriguez (2021) is the most relevant to our task. The paper focuses on sequential decision-making under uncertainty and generates counterfactual explanations that specify an alternative sequence of actions differing in at most $k$ actions from the observed sequence that could have led to a better outcome. However, by employing dynamic programming and iterating through all the possible states and actions, the method only solves tasks with low-dimensional states and action spaces, as later shown in the experiments. Additionally, only comparing which actions are changed is not meaningful in many applications because the entire paths are already different after taking different actions, even if the following actions are the same. Finally, this method also does not consider personalization to an agent when looking for a counterfactual sequence of actions. This method will be later used as a baseline for some simple tasks with low complexity.
@@ 2.3 Alternative Path Planning
In addition, there is prior work that find alternative policies for an agent by modifying its learning process (Racanière et al. 2017; Zahavy et al. 2021; Andrychowicz et al. 2017). However, those works aim to discover either more diverse or potentially better policies. While their ideas can help improve the agent’s training process, none of them could help solve finding an agent’s personalized recourse path.
@@ 3.1 Background and Notations
An MDP is a tuple M $\triangleq(\mathcal{S},\mathcal{A},P,R,\lambda$), where $\mathcal{S}=\{\mathbf{s}_{1},...,\mathbf{s}_{|\mathcal{S}|}\}$ and $\mathcal{A}=\{a_{1},...,a_{|\mathcal{A}|}\}$ are the set of states and actions, respectively. Here, $|\mathcal{S}|$, $|\mathcal{A}|$ are the cardinalities of set $\mathcal{S}$ and $\mathcal{A}$ respectively. At time $t$, the agent at state $\mathbf{s}_{t}$ takes an action $a_{t}$ and moves to state $\mathbf{s}_{t+1}$ with a transition probability $P(\mathbf{s}_{t+1}|\mathbf{s}_{t},a_{t})$ and receives a reward, $r_{t}$, which is determined by a reward function $R(\mathbf{s}_{t},a_{t})$. Behavior of the agent is defined by a policy function $\mathcal{P}_{A}(\mathbf{s}_{t},a_{t})$, which defines the probability that the agent $A$ will choose action $a_{t}$ given state $\mathbf{s}_{t}$. We denote $\tau=<\mathbf{s}_{1},\mathbf{s}_{2},...,\mathbf{s}_{T}>$ as a path, which is a sequence of states.
@@
There are many algorithms for training a reinforcement learning agent. In this paper, we employ deep Q-networks (DQN) to train a recourse policy. In a DQN algorithm, the agent explores the environment over multiple episodes, where each episode consists of a sequence of interactions with the environment from start to finish. For each episode, at each time step $t$, we record the agent’s current state $\mathbf{s}_{t}$, action $a_{t}$, the next state $\mathbf{s}_{t+1}$ and the reward $r_{t}$. Those tuples $(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})$ are stored in an experience replay $B$. DQN trains a deep neural network parameterized by $\theta$, denoted as $Q_{\theta}(\mathbf{s}_{t},a_{t})$, to help the agent maximize its cumulative reward from the environment. To stabilize the training, a target network $Q_{\theta^{\prime}}(\mathbf{s}_{t},a_{t})$ is introduced, which is parameterized by $\theta^{\prime}$ and updated periodically by setting $\theta^{\prime}=\theta$. This delay in updating the target network enhances the stability of the original network $Q_{\theta}$. With the experience replay buffer and the target network in place, the DQN algorithm minimizes the following loss function:
$\begin{split}L(\theta)=\operatorname{\mathbb{E}}_{(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})\sim\mathcal{V}(B)}[(r_{t}+\\ \gamma*\max_{a_{t}}Q_{\theta^{\prime}}(\mathbf{s}_{t+1},a_{t})-Q_{\theta}(\mathbf{s}_{t},a_{t}))^{2}],\end{split}$ (1)
where $\lambda\in$ [0, 1) is the discount factor. Mini batches of tuples $(\mathbf{s}_{t},a_{t},\mathbf{s}_{t+1},r_{t})$ are sampled from the replay buffer $B$ with sample strategy $\mathcal{V}$, such as uniform distribution. Note that other extensions from DQN such as prioritized replay buffer (Schaul et al. 2015) or dueling networks (Wang et al. 2016) can also be applied.
@@ 3.2 Problem Formulation
Given an agent $A$ characterized by its policy function $\mathcal{P}_{A}(\mathbf{s},a)$ and an original path $\tau_{0}=<\mathbf{s}^{\tau_{0}}_{1},\mathbf{s}^{\tau_{0}}_{2},...,\mathbf{s}^{\tau_{0}}_{T_{0}}>$ made by $A$, our goal is to find a different path $\tau_{r}=<\mathbf{s}^{\tau_{r}}_{1},\mathbf{s}^{\tau_{r}}_{2},...,\mathbf{s}^{\tau_{r}}_{T_{r}}>$ of sequence $\tau_{0}$ that achieves certain goal $\mathcal{G}$ while being personalized to agent $A$, at both path level and policy level.
Goal The goal $\mathcal{G}$ could represent different constraints or desired outcomes, depending on the task. For example, it could require agent $A$ to arrive at a particular destination state $\mathbf{s}_{d}$; or it could require the overall reward to be larger than some pre-defined threshold (e.g., a driver reaching the final destination faster); or in the context of sequence data in a supervised learning setting, it could require the sequence to be classified as a particular class. The goal can be formulated as a continuous outcome, the higher (lower) the better, or a binary outcome, whether certain criteria are met.
Personalization at Path Level Path-level personalization requires the $\tau_{r}$ to be as similar to the original path $\tau_{0}$ as possible. For example, the new path $\tau_{r}$ recommended to the driver should have many overlapped states with the original path $\tau_{0}$ picked by the driver.
Measuring the similarity (distance) between two sequences, $\tau_{0}$ and $\tau_{r}$, is a non-trivial task. The problem becomes even more challenging if the sequences have different lengths. For this reason, traditional metrics such as Euclidean or Manhattan distance metrics do not apply. In our work, we measure the distance between $\tau_{0}$ and any path $\tau_{r}$ by Levenshtein distance (Levenshtein et al. 1966), i.e., the edit distance, denoted as $d(\tau_{0},\tau_{r})$, which counts the minimum number of operations (insertion, deletion, and substitution ) to convert $\tau_{0}$ to $\tau_{r}$. For instance, the edit distance between the sequences ”LRLRL” and ”LRLRR” is 1 (only the last character differs), while the distance between ”LRLRL” and ”LRRLLR” is 3 (three characters need to be changed to make them identical).
Personalization at Policy Level This property requires the new path $\tau_{r}$ to align closely with the behavior of agent $A$. For example, if $\tau_{r}$ is a new path recommended to a driver aiming to reach a destination quickly, we might suggest a highway for an experienced driver and a local path for a novice driver. Essentially, the recommended path $\tau_{r}$ should respect agent $A$’s inherent behavior patterns.